OpenStudy (anonymous):

$5,000 is invested at 3% interest. How much money must be invested at 5% interest so that the total interest from the two investments is $275 after one year?

OpenStudy (shinalcantara):

let x be the amount invested @ 5% You'll have the equation, 5,000(.03)+x(.05)=275 .05x=275-5,000(.03) x=2,500